I have a question: how do u solve the ones below?
1) x^2+7x+10
2) 2x^2-50

Answer by tutorcecilia(2144)

1) x^2+7x+10 [Factor if possible]
(x+2)(x+5) [Set each factor equal to zero]
(x+2) = 0
(x+5) = 0
.
x+2=0 x=-2 [Solve for x]
.
x+5=0
x=-5 [Solve for x]
.
x = -2 or x=-5 [Final answers]